The solubility of N2 (g) at 25 C and 1.250 atm N2 (g) pressure is 0.02136 N2 per liter of water. If a sample of N2 (g) under a pressure of 12.15 atm is kept in contact with 3.852 L of water at 25.0 C, what mass (in grams) of N2 will dissolve?
1
Expert's answer
2019-09-17T08:26:32-0400

Henry's Law decides the solubility of gases in liquids.

S2=S1*P2/P1 = (0.02136*12.15)/(3.852*1.25) = 0.053 g